# FedAPTA: Federated Multi-task Learning in Computing Power Networks with Adaptive Layer-wise Pruning and Task-aware Aggregation

### v1

@@ Abstract
Federated Learning (FL) has shown considerable promise in Computing Power Networks (CPNs) for privacy protection, efficient data utilization, and dynamic collaboration. Although it offers practical benefits, applying FL in CPNs continues to encounter a major obstacle, i.e., multi-task deployment. However, existing work mainly focuses on mitigating FL’s computation and communication overhead of a single task while overlooking the computing resource wastage issue of heterogeneous devices across multiple tasks in FL under CPNs. To tackle this, we design FedAPTA, a federated multi-task learning framework in CPNs. FedAPTA alleviates computing resource wastage through the developed layer-wise model pruning technique, which reduces local model size while considering both data and device heterogeneity. To aggregate structurally heterogeneous local models of different tasks, we introduce a heterogeneous model recovery strategy and a task-aware model aggregation method that enables the aggregation through infilling local model architecture with the shared global model and clustering local models according to their specific tasks. We deploy FedAPTA on a realistic FL platform and benchmark it against nine SOTA FL methods. The experimental outcomes demonstrate that the proposed FedAPTA considerably outperforms the state-of-the-art FL methods by up to 4.23%. Our code is available at https://github.com/Zhenzovo/FedCPN.
Index Terms:
@@
Computing Power Networks (CPNs), federated learning, multiple tasks, model pruning, heterogeneous aggregation
I Introduction
@@
In the past ten years, the proliferation of billions of Internet of Things (IoT) devices has led to an unprecedented increase in data generation [1]. In the context of Computing Power Networks (CPNs) [2], Federated Learning (FL) [3] has emerged as a promising solution for enabling distributed model training across heterogeneous computing devices while preserving data privacy [4, 5, 6, 7]. FL allows devices (i.e., participants of FL in CPNs) to locally train models based on their confidential data and regularly transmit the model updates to the central server for aggregation, which facilitates data processing on distributed computing devices where data is generated. Additionally, FL distributes model training across devices to avoid the transmission of raw data to the central server, thereby reducing communication overhead and preserving user privacy [3, 8].
@@
Although FL has been effectively applied in CPNs, FL still faces serious challenges when handling multi-task scenarios, particularly in terms of model deployment and aggregation. For model deployment, generally speaking, different tasks have different requirements for model expressiveness. Therefore, deploying the same model on devices that may perform different tasks can lead to computing resource wastage. For instance, a small model is usually sufficient for a device to perform a simple task, while a large model is normally required for a device to conduct a complex task. However, if all devices of a CPN are assigned the same large model regardless of their task complexity, computing resources of those performing simple tasks could be wasted. For model aggregation, if the central server aggregates local models of all devices directly without task differentiation, the resulting global model will fail to accurately learn the task-specific knowledge associated with each device, which can cause performance reduction of the global model.
@@
A number of pioneering studies have made significant efforts to tackle these two challenges. For the former, to prevent the waste of computing resources, task-specific model architecture should be generated. Some studies [9, 10, 11, 12, 13, 14] have explored model pruning strategies to address this issue. For example, FedMP [12] leverages a multi-armed bandit strategy to dynamically adjust the pruning ratio for each device, effectively reducing resource consumption while maintaining promising accuracy. With a similar goal of reducing resource consumption, FedLPS [13] proposes an adaptive channel-wise pruning method to generate lightweight local models, aiming to accommodate clients with varying computational capabilities in heterogeneous FL scenarios. However, existing research primarily considers device computational capabilities when designing pruning strategies, while overlooking the fundamental relationship between task requirements and model sizes. Besides, during the pruning process, these methods usually adopt a uniform pruning ratio for all layers to be pruned, without taking into account the varying importance levels of different layers, which can lead to poor performance of the pruned model. For the latter, to alleviate the issue of suboptimal global models caused by directly aggregating local models from different tasks, some studies cluster local models that belong to the same task before aggregation. For example, Clustered Federated Learning (CFL) [15] clusters devices based on the similarity of their gradient updates, enabling the learning of more specialized models within each cluster.
@@
In this paper, we propose a novel FEDerated multi-task learning framework with Adaptive layer-wise Pruning and Task-aware Aggregation (FedAPTA) for addressing the model deployment and aggregation challenges of multi-task learning in CPNs. Specifically, we allow each device to train a sub-model that is meticulously pruned to match its own local task requirements. Different from existing pruning methods [9, 10, 11], we assign a unique pruning ratio to each layer rather than applying a uniform pruning ratio for all layers of a model to reduce the computing resources wastage in multi-task scenarios in CPNs and maintain satisfactory training performance even under relatively high pruning ratios. To aggregate the heterogeneous pruned models, we introduce a heterogeneous model recovery algorithm. Unlike existing algorithms [11, 12] that only aggregate shared parameters across pruned local models, FedAPTA leverages the information from the latest global model to assist in the reconstruction of a consistent architecture from the pruned models. It mitigates the issue wherein parameters pruned by other devices are no longer able to contribute to the learning process. Additionally, since aggregating local models belonging to different tasks on the central server may lead to a suboptimal global model, we cluster devices that are performing the same task by their local models before aggregation.
@@
    We propose an adaptive layer-wise model pruning method for FedAPTA that innovatively allocates pruning ratios based on layer importance. This reduces the parameter size of local models while maintaining accuracy, allowing devices to learn heterogeneous models tailored to their task requirements and minimizing resource consumption in multi-task scenarios.
@@
    We integrate the developed FedAPTA into a realistic federated learning platform (FedML) and conduct comprehensive evaluations against nine SOTA frameworks. Experimental outcomes show that FedAPTA maintains superior model accuracy compared with the aforementioned FL frameworks and alleviates unnecessary resource consumption.
@@
TABLE I: List of key notations.
Notation Description
$C$ A set of devices
$i$ The $i_{th}$ device in $C$
$D_{i}$ The local dataset of device $i$
$\rho_{i}$ The pruning ratio for the model of device $i$
$R$ Total FL rounds
$r$ The $r_{th}$ FL round
$w$ The initialized global model weights
$w_{i}$ The model weights of device $i$
$\hat{w_{i}}$ The pruned model weights of device $i$
$M_{i}$ The pruning mask matrix of device $i$
$\mathbb{R}^{n}$ The n-dimensional real vector space
$P_{i}$ A set of layer-wise pruning ratios for the model of device $i$
$\rho_{k}$ The pruning ratio for layer $k$
$k$ The $k_{th}$ trainable convolutional layer of a local model
$I$ A set of importance scores for each candidate layer
$I_{k}$ The importance score of layer $k$
$N$ A set of parameter numbers of each candidate layer
$N_{k}$ The parameter count of layer $k$
$d_{cos}$ Cosine distance
$\Delta w_{i}$ The local model updates of device $i$
$T$ A set of tasks
$t$ The $t_{th}$ task in $T$
$D_{t}$ The set of local datasets on devices belonging to task $t$
$w_{t}$ The global model for task $t$
II Methodology
II-A Overview
In this study, we propose FedAPTA, a FL framework specifically designed to efficiently train task-specific models across heterogeneous devices in CPNs, where different devices may undertake different tasks. Unlike traditional FL frameworks, which focus on training and maintaining a shared model structure for all devices, FedAPTA allows each device to possess a local model customized for its task requirements. The overall operational workflow of FedAPTA is illustrated in Fig. 1, which includes two fundamental steps: the training of local models on devices and the aggregation of models on the central server. Particularly, the central server uses a pre-trained model as the initialized global model $w$ and distributes it to all devices. $w$ can come from pre-trained models trained on openly accessible datasets. To facilitate efficient training under multi-task scenarios in CPNs, we introduce an adaptive layer-wise model pruning method, which will be described in detail in Section II-B. In addition, regarding to the model aggregation process, there are two central challenges that must be addressed: on one hand, to aggregate the pruned heterogeneous models, we utilize information from the latest global model to restore the complete structure of the local models, facilitating the aggregation of the pruned models (as described in Section II-C); on the other hand, to aggregate different task-specific models, we cluster all devices by their local models to distinguish their tasks (as elaborated in Section II-D). The complete implementation process of the proposed method is presented in Algorithm 1, and the key symbols used in this paper are listed in Table I.
We outline the steps of FL as follows: 1) The central server initializes the global model $w$ and distributes it to all devices. 2) Each device $i$ determines an appropriate pruning ratio $\rho_{i}$ based on its individual task requirements, and applies the adaptive layer-wise model pruning technique to obtain a customized local model $\hat{w_{i}}$. 3) Each device $i$ trains its local model using its private dataset $D_{i}$, and subsequently uploads the pruned local model $\hat{w_{i}}$ and the mask matrix $M_{i}$ to the central server. 4) The central server restores the pruned parts of each local model $\hat{w_{i}}$ using the corresponding mask matrix $M_{i}$ and the latest global model $w_{t}$, where $t$ is the task executed by device $i$, thereby restoring local models to a consistent structural format suitable for aggregation. 5) The central server conducts clustering over the reconstructed local models in order to distinguish models based on their respective tasks. 6) The central server aggregates models within each identified cluster, thereby aggregating local models related to the same task $t$. 7) The central server distributes the aggregated task-specific global models $w_{t}$ to the corresponding devices for further training.
II-B Adaptive Layer-wise Model Pruning
@@
Existing FL frameworks, such as FedAvg [3], commonly adhere to a relatively fixed scheme: devices are assigned local models that share the same architecture as the global model, while local training and updates are conducted on this unified model architecture. While this design simplifies the model aggregation and synchronization processes, it exhibits significant limitations in practice. Specifically, due to varying tasks across devices in CPNs, enforcing the optimization of an identical model structure across all devices often results in computing resources wastage, thereby restricting the applicability of FL in device-heterogeneous CPN environments.
@@
To address these challenges, we propose FedAPTA. Distinct from existing FL frameworks that require all devices to train an identical model, in FedAPTA, after the global model is sent from the central server, each device $i\in C$ prunes the global model using the proposed adaptive layer-wise model pruning method in order to derive a local model that is compatible with its own task requirements. In this context, the pruning ratio $\rho_{i}$ is determined by device $i$, based on an assessment of its task requirements. For example, devices engaged in relatively simple tasks tend to adopt higher pruning ratios. Through this adaptive layer-wise pruning process, each device is capable of learning a local model that is structurally sparse. This pruning strategy enables devices with different task requirements to participate in FL training using models that are tailored to their needs, thereby avoiding unnecessary computing resource consumption.
@@
During the FL training process, after receiving the global model, each device $i$ performs a pruning operation on the model using a pruning ratio $\rho_{i}$. In contrast to approaches such as FedDrop [9], which adopts a uniform pruning ratio across all participating devices in FL, FedAPTA allows each device to decide its pruning ratio $\rho_{i}$ based on its unique task requirements, improving FL flexibility. What’s more, for each trainable convolutional layer of the local model, FedAPTA applies a differentiated pruning ratio that is adaptively assigned according to the assessed importance of each candidate layer to better preserve critical features and improve the balance between pruning effectiveness and model performance.
@@
To begin with, for the model deployed on each device $i$, FedAPTA evaluates the importance score of each candidate layer $k$ within the model $w_{i}$ by employing the L1 norm [16], obtaining the importance score vector $I=[I_{1},\cdots,I_{k},\cdots,I_{n}]$. Subsequently, while ensuring that the overall pruning ratio of the model remains consistent with the predefined whole pruning ratio target $\rho_{i}$, the pruning ratio $\rho_{k}$ for each candidate layer $k$ is assigned based on its parameter quantity and importance score. Specifically, let the parameter count of each candidate layer $k$ be represented by the vector $N=[N_{1},\cdots,N_{k},\cdots,N_{n}]$; this vector is then reordered according to the descending order of $I$. The objective is to assign a pruning ratio $\rho_{k}$ to each candidate layer $k$, where $\rho_{k}$ is between 0 and 1, such that the aggregate pruning across all layers adheres to the specified overall pruning ratio $\rho_{i}$, while the sequence of assigned pruning ratios exhibits a non-decreasing structure, it means that the layers that are deemed less important are assigned a pruning ratio that is no less than the ratios of more important layers. Let $P_{i}=[\rho_{1},\cdots,\rho_{k},\cdots,\rho_{n}]$ represent the optimization variable, then the formulation of the optimization problem is given as follows:
$\min_{P_{i}\in\mathbb{R}^{n}}$ 0 0,0 , (1)
@@
This method enables heterogeneity and sparsity of model structures, effectively adapting to the different task requirements of devices, thereby reducing computing resources wastage and enhancing the practical applicability of FL in CPNs.
@@
II-C Heterogeneous Model Recovery
Adaptive layer-wise model pruning offers a significant advantage in FL, particularly when applied across devices with diverse tasks in CPNs. By systematically removing less important parameters in a model on a per-layer basis, this method effectively reduces the overall size of the local models. Consequently, it significantly reduces computational cost, which is especially beneficial for devices with different tasks. This approach allows each device to determine the pruning ratio for its local model according to its task requirements, enabling more flexible and on-demand pruning. By enabling devices with different tasks to participate in FL using models with different sizes, it reduces computing resources wastage caused by deploying a uniform model architecture on heterogeneous devices in multi-task scenarios.
However, despite these benefits, the proposed adaptive layer-wise pruning technique inherently includes structural heterogeneity among devices’ models. Specifically, since each device independently prunes its local model according to its task requirements, the resulting model architectures can vary significantly from one device to another. It presents a major challenge for model aggregation, which is a fundamental operation in FL, thereby making models incompatible with popular and widely adopted model aggregation algorithms, such as FedAvg [3], as mismatched architectures prevent straightforward averaging or fusion. As a result, application of such algorithms to models with heterogeneous structures may result in degraded global model performance or even complete aggregation failure. To cope with these structural discrepancies, several heterogeneous model aggregation approaches have been proposed recently. For example, FedMP [12] focuses on aggregating the parameters that are shared among heterogeneous local models. Nevertheless, if certain parameters of a device’s local model have been pruned on other devices, the device will be unable to benefit from those parameters.

### v2

# FedAPTA: Federated Multi-task Learning for Heterogeneous Devices with Adaptive Layer-wise Pruning and Task-aware Aggregation

@@ Abstract
Federated Learning (FL) has shown considerable promise in Machine Learning (ML) across numerous devices for privacy protection, efficient data utilization, and dynamic collaboration. However, mobile devices typically have limited and heterogeneous computational capabilities, and different devices may even have different tasks. This client heterogeneity is a major bottleneck hindering the practical application of FL. Existing work mainly focuses on mitigating FL’s computation and communication overhead of a single task while overlooking the computing resource heterogeneity issue of different devices in FL. To tackle this, we design FedAPTA, a federated multi-task learning framework. FedAPTA overcomes computing resource heterogeneity through the developed layer-wise model pruning technique, which reduces local model size while considering both data and device heterogeneity. To aggregate structurally heterogeneous local models of different tasks, we introduce a heterogeneous model recovery strategy and a task-aware model aggregation method that enables the aggregation through infilling local model architecture with the shared global model and clustering local models according to their specific tasks. We deploy FedAPTA on a realistic FL platform and benchmark it against nine SOTA FL methods. The experimental outcomes demonstrate that the proposed FedAPTA considerably outperforms the state-of-the-art FL methods by up to 4.23%. Our code is available at https://github.com/Zhenzovo/FedAPTA.
keywords:
@@
Federated learning , Client heterogeneity , Multiple tasks , Model pruning , Heterogeneous aggregation
journal: Nuclear Physics B
[…]
[a]organization=School of Future Science and Engineering,addressline=Soochow University, city=Suzhou, state=Jiangsu, country=China
[…]
[c]organization=Key Laboratory of General Artificial Intelligence and Large Models in Provincial Universities (Soochow University),city=Suzhou, state=Jiangsu, country=China
[…]
[b]organization=China Mobile (Suzhou) Software Technology Co., Ltd.,city=Suzhou, state=Jiangsu, country=China
1 Introduction
@@
In the past ten years, the proliferation of billions of Internet of Things (IoT) devices has led to an unprecedented increase in data generation [khan2021federated_khaniot]. Federated Learning (FL) [mcmahan2017communicationfedavg] has emerged as a promising solution for enabling distributed model training across heterogeneous devices while preserving data privacy [sun2022fedtar, guendouzi2023systematic, yuan2023computing_split, wei2024fedact, lin2024game_cpngame, wang2025novel_hete]. FL allows devices (i.e., participants of FL) to locally train models based on their confidential data and regularly transmit the model updates to the central server for aggregation, which facilitates data processing on distributed computing devices where data is generated. Additionally, FL distributes model training across devices to avoid the transmission of raw data to the central server, thereby reducing communication overhead and preserving user privacy [mcmahan2017communicationfedavg, ezzeldin2023fairfed].
@@
Although FL has been effectively applied, it still faces serious challenges when handling multi-task scenarios with heterogeneous devices, particularly in terms of model deployment and aggregation. For model deployment, in most cases, heterogeneous devices participating in FL exhibit significant differences in computational capability, and many of them have limited resources. Therefore, if the same model is deployed on all devices without any adaptation, those with weaker computational capability will be unable to properly train the model or effectively participate in the FL process. This may hinder the progress of global model training and even cause the entire FL procedure to stall, ultimately affecting the normal training of the global model. For model aggregation, if the central server aggregates local models of all devices directly without task differentiation, the resulting global model will fail to accurately learn the task-specific knowledge associated with each device, which can cause performance reduction of the global model.
@@
A number of pioneering studies have made significant efforts to tackle these two challenges. For the former, to enable devices with weaker computational capabilities to properly participate in the FL process, model architectures should be generated according to their computing resources. Some studies [caldas2018expanding_feddrop, alam2022fedrolex, li2021hermes, jiang2022fedmp, jia2024fedlps, 11052838] have explored model pruning strategies to address this issue. For example, FedMP [jiang2022fedmp] leverages a multi-armed bandit strategy to dynamically adjust the pruning ratio for each device, effectively adapting to heterogeneous devices while maintaining promising accuracy. For the same goal of accommodating the heterogeneous computational capabilities of devices, FedLPS [jia2024fedlps] proposes an adaptive channel-wise pruning method to generate lightweight local models. However, during the pruning process, existing research usually adopt a uniform pruning ratio for all layers to be pruned, without taking into account the varying importance levels of different layers, which can lead to poor performance of the pruned model. For the latter, to alleviate the issue of suboptimal global models caused by directly aggregating local models from different tasks, some studies cluster local models that belong to the same task before aggregation. For example, Clustered Federated Learning (CFL) [sattler2020clustered_cfl] clusters devices based on the similarity of their gradient updates, enabling the learning of more specialized models within each cluster.
@@
In this paper, we propose a novel FEDerated multi-task learning framework for heterogeneous devices with Adaptive layer-wise Pruning and Task-aware Aggregation (FedAPTA) for addressing the model deployment and aggregation challenges in FL. Specifically, we allow each device to train a sub-model that is meticulously pruned to match its own local computing resources. Different from existing pruning methods [caldas2018expanding_feddrop, alam2022fedrolex, li2021hermes], we assign a unique pruning ratio to each layer rather than applying a uniform pruning ratio for all layers of a model to maintain satisfactory training performance even under relatively high pruning ratios. To aggregate the heterogeneous pruned models, we introduce a heterogeneous model recovery algorithm. Unlike existing algorithms [li2021hermes, jiang2022fedmp] that only aggregate shared parameters across pruned local models, FedAPTA leverages the information from the latest global model to assist in the reconstruction of a consistent architecture from the pruned models. It mitigates the issue wherein parameters pruned by other devices are no longer able to contribute to the learning process. Additionally, since aggregating local models belonging to different tasks on the central server may lead to a suboptimal global model, we cluster devices that are performing the same task by their local models before aggregation.
@@
    We propose an adaptive layer-wise model pruning method for FedAPTA that innovatively allocates pruning ratios based on layer importance. This reduces the parameter size of local models while maintaining accuracy, allowing devices to learn heterogeneous models tailored to their computational capabilities and ensuring the efficient progress of FL.
@@
    We integrate the developed FedAPTA into a realistic federated learning platform (FedML) and conduct comprehensive evaluations against nine SOTA frameworks. Experimental outcomes show that FedAPTA maintains superior model accuracy compared with the aforementioned FL frameworks.
Table 1: List of key notations.
Notation Description
$C$ A set of devices
$i$ The $i_{th}$ device in $C$
$D_{i}$ The local dataset of device $i$
$\rho_{i}$ The pruning ratio for the model of device $i$
$R$ Total FL rounds
$r$ The $r_{th}$ FL round
$w$ The initialized global model weights
$w_{i}$ The model weights of device $i$
$\hat{w_{i}}$ The pruned model weights of device $i$
$M_{i}$ The pruning mask matrix of device $i$
$\mathbb{R}^{n}$ The n-dimensional real vector space
$P_{i}$ A set of layer-wise pruning ratios for the model of device $i$
$\rho_{k}$ The pruning ratio for layer $k$
$k$ The $k_{th}$ trainable convolutional layer of a local model
II A set of importance scores for each candidate layer
$I_{k}$ The importance score of layer $k$
NN A set of parameter numbers of each candidate layer
$N_{k}$ The parameter count of layer $k$
$d_{cos}$ Cosine distance
$\Delta w_{i}$ The local model updates of device $i$
$T$ A set of tasks
$t$ The $t_{th}$ task in $T$
$D_{t}$ The set of local datasets on devices belonging to task $t$
$w_{t}$ The global model for task $t$
2 Methodology
2.1 Overview
In this study, we propose FedAPTA, a FL framework specifically designed to efficiently train task-specific models across heterogeneous devices, where different devices have different computing resources and may undertake different tasks. Unlike traditional FL frameworks, which focus on training and maintaining a shared model structure for all devices, FedAPTA allows each device to possess a local model customized for its computational capability. The overall operational workflow of FedAPTA is illustrated in Fig. 1, which includes two fundamental steps: the training of local models on devices and the aggregation of models on the central server. Particularly, the central server uses a pre-trained model as the initialized global model $w$ and distributes it to all devices. $w$ can come from pre-trained models trained on openly accessible datasets. To facilitate efficient training in scenarios with heterogeneous client computational capabilities, we introduce an adaptive layer-wise model pruning method, which will be described in detail in Section $2.$ In addition, regarding to the model aggregation process, there are two central challenges that must be addressed: on one hand, to aggregate the pruned heterogeneous models, we utilize information from the latest global model to restore the complete structure of the local models, facilitating the aggregation of the pruned models (as described in Section 2.3); on the other hand, to aggregate different task-specific models, we cluster all devices by their local models to distinguish their tasks (as elaborated in Section 2.4). The complete implementation process of the proposed method is presented in Algorithm 1, and the key symbols used in this paper are listed in Table 1.
We outline the steps of FL as follows: 1) The central server initializes the global model $w$ and distributes it to all devices. 2) Each device $i$ determines an appropriate pruning ratio $\rho_{i}$ based on its individual computational capability, and applies the adaptive layer-wise model pruning technique to obtain a customized local model $\hat{w_{i}}$. 3) Each device $i$ trains its local model using its private dataset $D_{i}$, and subsequently uploads the pruned local model $\hat{w_{i}}$ and the mask matrix $M_{i}$ to the central server. 4) The central server restores the pruned parts of each local model $\hat{w_{i}}$ using the corresponding mask matrix $M_{i}$ and the latest global model $w_{t}$, where $t$ is the task executed by device $i$, thereby restoring local models to a consistent structural format suitable for aggregation. 5) The central server conducts clustering over the reconstructed local models in order to distinguish models based on their respective tasks. 6) The central server aggregates models within each identified cluster, thereby aggregating local models related to the same task $t$. 7) The central server distributes the aggregated task-specific global models $w_{t}$ to the corresponding devices for further training.
@@ 2.1 Overview
2.2 Adaptive Layer-wise Model Pruning
@@
Existing FL frameworks, such as FedAvg [mcmahan2017communicationfedavg], commonly adhere to a relatively fixed scheme: devices are assigned local models that share the same architecture as the global model, while local training and updates are conducted on this unified model architecture. While this design simplifies the model aggregation and synchronization processes, it exhibits significant limitations in practice. Specifically, since devices in FL have different computational capabilities, enforcing the optimization of an identical model structure across all devices often leads to inefficiency, thereby restricting the applicability of FL in heterogeneous devices environments.
@@
To address these challenges, we propose FedAPTA. Distinct from existing FL frameworks that require all devices to train an identical model, in FedAPTA, after the global model is sent from the central server, each device $i\in C$ prunes the global model using the proposed adaptive layer-wise model pruning method in order to derive a local model that is compatible with its computational capability. In this context, the pruning ratio $\rho_{i}$ is determined by device $i$, based on an assessment of its computing resources. Through this adaptive layer-wise pruning process, each device is capable of learning a local model that is structurally sparse. This pruning strategy enables heterogeneous devices to train models that are tailored to their computational capabilities, thereby avoiding the impact on the efficiency of FL.
@@
During the FL training process, after receiving the global model, each device $i$ performs a pruning operation on the model using a pruning ratio $\rho_{i}$. In contrast to approaches such as FedDrop [caldas2018expanding_feddrop], which adopts a uniform pruning ratio across all participating devices in FL, FedAPTA allows each device to decide its pruning ratio $\rho_{i}$ based on its unique computational capability, improving the flexibility of FL. What’s more, for each trainable convolutional layer of the local model, FedAPTA applies a differentiated pruning ratio that is adaptively assigned according to the assessed importance of each candidate layer to better preserve critical features and improve the balance between pruning effectiveness and model performance.
@@
To begin with, for the model deployed on each device $i$, FedAPTA evaluates the importance score of each candidate layer $k$ within the model $w_{i}$ by employing the L1 norm [liu2017learning_pr1], obtaining the importance score vector $I=[I_{1},\cdots,I_{k},\cdots,I_{n}]$. Subsequently, while ensuring that the overall pruning ratio of the model remains consistent with the predefined whole pruning ratio target $\rho_{i}$, the pruning ratio $\rho_{k}$ for each candidate layer $k$ is assigned based on its parameter quantity and importance score. Specifically, let the parameter count of each candidate layer $k$ be represented by the vector $N=[N_{1},\cdots,N_{k},\cdots,N_{n}]$; this vector is then reordered according to the descending order of II. The objective is to assign a pruning ratio $\rho_{k}$ to each candidate layer $k$, where $\rho_{k}$ is between 0 and 1, such that the aggregate pruning across all layers adheres to the specified overall pruning ratio $\rho_{i}$, while the sequence of assigned pruning ratios exhibits a non-decreasing structure, it means that the layers that are deemed less important are assigned a pruning ratio that is no less than the ratios of more important layers. Let $P_{i}=[\rho_{1},\cdots,\rho_{k},\cdots,\rho_{n}]$ represent the optimization variable, then the formulation of the optimization problem is given as follows:
$\min_{P_{i}\in\mathbb{R}^{n}}$ $0,$ (1)
@@
This method enables heterogeneity and sparsity of model structures, effectively adapting to the different computational capabilities of devices, thereby enhancing the practical applicability of FL.
@@
2.3 Heterogeneous Model Recovery
Adaptive layer-wise model pruning offers a significant advantage in FL, particularly when applied across devices with different computing resources. By systematically removing less important parameters in a model on a per-layer basis, this method effectively reduces the overall size of the local models. Consequently, it significantly reduces computational cost, which is especially beneficial for heterogeneous devices. This approach allows each device to determine the pruning ratio for its local model according to its computational capability, enabling more flexible and on-demand pruning. By enabling heterogeneous devices to participate in FL using models with different sizes, it avoids a decrease in the efficiency of FL caused by deploying a uniform model architecture on devices with heterogeneous computing resources.
However, despite these benefits, the proposed adaptive layer-wise pruning technique inherently includes structural heterogeneity among devices’ models. Specifically, since each device independently prunes its local model according to its computational capability, the resulting model architectures can vary significantly from one device to another. It presents a major challenge for model aggregation, which is a fundamental operation in FL, thereby making models incompatible with popular and widely adopted model aggregation algorithms, such as FedAvg [mcmahan2017communicationfedavg], as mismatched architectures prevent straightforward averaging or fusion. As a result, application of such algorithms to models with heterogeneous structures may result in degraded global model performance or even complete aggregation failure. To cope with these structural discrepancies, several heterogeneous model aggregation approaches have been proposed recently. For example, FedMP [jiang2022fedmp] focuses on aggregating the parameters that are shared among heterogeneous local models. Nevertheless, if certain parameters of a device’s local model have been pruned on other devices, the device will be unable to benefit from those parameters.